OpenStudy (anonymous):

A particle which is moving in a straight line with constant acceleration covers distances of 10m and 15m in two successive seconds. Find the acceleration

OpenStudy (anonymous):

@TuringTest @UnkleRhaukus @precal

OpenStudy (anonymous):

i ve just started learning this at my class...but i know those:)

OpenStudy (precal):

derivative and calculus correct?

OpenStudy (foolaroundmath):

I'll assume you are familiar with the equations of motion, particularly this one \[s = ut + \frac{1}{2}at^{2}\] Let us take the starting time to be t. In time t -> t+1, it covers 10m and from time t+1 -> t+2 it covers 15m Now instead of using t+1 -> t+2, i'll make use of the fact that it covered 25m from t->t+2 This is used, because in both cases the initial velocity 'u' is the same. 10 = u(t+1) + 0.5a(t+1)^2 - [ut + 0.5at^2] 25 = u(t+2) + 0.5a(t+2)^2 - [ut + 0.5at^2] Can you take it from here? The answer is still a couple steps away

OpenStudy (precal):

listen to FoolAroundMath

OpenStudy (anonymous):

define two successive seconds?

OpenStudy (foolaroundmath):

t->t+1 is the first second, t+1->t+2 is the second successive second

OpenStudy (unklerhaukus):

\[\vec a=\frac{\vec v_1+\vec v_2}{2}=\frac{\frac{\Delta x_1}{\Delta t_1}+\frac{\Delta x_2}{\Delta t_2}}{2}=\frac{\frac{10\text m}{1\text s}+\frac{15\text m}{1\text s}}{2}\]

OpenStudy (anonymous):

@FoolAroundMath why did u take 0.5?

OpenStudy (foolaroundmath):

Isn't that the second equation of motion? \[s = ut + \frac{1}{2}at^{2}\] adding a bit more to my answer, The displacement in t->t+1 = total displacement from time 0 to t+1 - total displacement from time 0 to t. The displacement in t->t+2 = total displacement from time 0 to t+2 - total displacement from time 0 to t

OpenStudy (anonymous):

isnt there any other easy way to do this question?

OpenStudy (foolaroundmath):

Can't think of anything simpler at the moment. :(

OpenStudy (anonymous):

hm anyway thanks :)

OpenStudy (anonymous):

@ash2326 @TuringTest

OpenStudy (anonymous):

no...><

OpenStudy (maheshmeghwal9):

wait a minute I m seeing formula again:)

OpenStudy (maheshmeghwal9):

yeah i was right \[\overrightarrow a=\frac{\overrightarrow S_2 -\overrightarrow S_1}{t^2}\]

OpenStudy (maheshmeghwal9):

\[\overrightarrow a= \frac{15-10}{1^2}=5 m/s^2\]

OpenStudy (maheshmeghwal9):

@thushananth01 gt it?

OpenStudy (anonymous):

but how, i never learnt tht equation?

OpenStudy (maheshmeghwal9):

wait a minute

OpenStudy (maheshmeghwal9):

\[\overrightarrow a=\frac{\overrightarrow v_2-\overrightarrow v_1}{t}\]\[\implies \overrightarrow a = \frac{\frac {\overrightarrow S_2}{t}-\frac{\overrightarrow S_1}{t}}{t}\]\[\implies \overrightarrow a = \frac{\overrightarrow S_2-\overrightarrow S_1}{t^2}\]

OpenStudy (maheshmeghwal9):

gt it now?

OpenStudy (anonymous):

yes thnks :)

OpenStudy (maheshmeghwal9):

yw:) so correct answer is 5m/s^2

OpenStudy (anonymous):

yes

OpenStudy (maheshmeghwal9):

^_^

OpenStudy (unklerhaukus):

thats what i was trying to do

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

mahesh .. so .. why is the time 1 ... it should be 2